What are concentration cells? Emf of the cell Cd | CdSO4 (X M)||CdSO4 (0.025 M)| Cd at 28oC is 0.035 V. Find the Concentration of CdSO4 at anode. Given R=8.314 J/K/mol, F=96500 C.
1
Expert's answer
2019-09-13T05:50:02-0400

 Concentration cells are galvanic  cells made of two half-cells, each of which containing the same electrodes, but different concentration.


EMF of a concentration cell is given by


"E_{cell}=\frac{RT}{nF}ln\frac{[C_2]}{[C_1]}=\frac{0.0591}{n}log\frac{[C_2]}{[C_1]}"

"C_2=0.025\ M\\C_1=X\\E_{cell}=0.035\ V\\n=2
\\now,\\0.035=\frac{0.0591}{2}log\frac{0.035}{X}"


"1.1844=log\frac{0.035}{X}\\or\ \\\\\frac{0.035}{X}=antilog(1.1844)=15.29
\\"

"or\\X=\frac{0.035}{15.29}=0.002289\ M"


So, concentration of"\ \ CdSO_4" at anode = "0.002289\ M"